At an amusement park, a swimmer uses a water slide
Answer #1

Solution(3)

                               Swimmers\; start\; from \; rest\; means \; its \; initial \; speed \; is\; Zero\; (u=0)

                               vertical \; height\; is\; given\; (h)\; =2.31m

                              According \; to\; the\; question\; we\; need \; to\; calculate\; the\; final\; speed

                 Apply\; Newton's\; third \; law

                                                           v^{2}=u^{2}+2gh

                                                         v^{2}=0+2\times 9.8\times 2.31m

                                                           v=\sqrt{45.276}

                                                           \mathbf{v=6.72\; m/s}\; \; \; (Ans)

Solution(4)

                             The\; initial\; speed \; of \; the\; pebble\; is\; given \; (u)\; =\; 14 m/s

                             The\; height\; of\; building\; is\; (h)\; =31.0 m

                              And\; we\; need \; to\; calculate\; the\; final\; speed \; (v):

                                     Again\; apply\; Newton's \; third \; motion\; law:

                                                     v^{2}=u^{2}+2gh

                                                      v^{2}=(14)^{2}+2\times 9.8\times 31

                                                        \mathbf{v^{2}=803.6 m/s}

                                                        \mathbf{v=28.34 m/s}\; \; \; \; (Ans)

Solution(5)      Now\; here\; apply\; energy\; conservation:

                                  \frac{1}{2}mv_{i}^{2}=\frac{1}{2}mv_{f}^{2}+mgh

                                 \frac{1}{2}v_{i}^{2}=\frac{1}{2}v_{f}^{2}+gh

                                 \frac{1}{2}(52)^{2}=\frac{1}{2}(47.1)^{2}+9.8\times h

                                 1352=1109.2+9.8\times h

                                9.8\times h=242.8

                                \mathbf{ h_{Max}=24.77\; m}\; \; \; (Ans)